# An Introduction to Reinforcement Learning: Fundamental Concepts and Practical Applications

### v1

@@ 6.3.2 Off-Policy Methods
Off-policy methods involve learning the value of the optimal policy independently of the agent’s actions. In these methods, we distinguish between two types of policies: the behavior policy and the target policy that are discussed later in this chapter. The behavior policy explores the environment, while the target policy aims to improve performance based on the gathered experience. This allows for a more exploratory behavior policy while learning an optimal target policy. A significant advantage of off-policy methods is that they can learn from data generated by any policy, not just the one currently being followed, making them highly flexible and sample-efficient.
Example: Consider a recommendation system for an online streaming service like Netflix. The behavior policy in this system could be a strategy that recommends a wide variety of content to users, ensuring that the system explores different genres, new releases, and less popular titles. This exploration helps gather diverse data about user preferences and content performance. Simultaneously, the target policy aims to optimize recommendations to maximize user engagement and satisfaction. It learns from the data generated by the behavior policy, identifying patterns and preferences to recommend content that users are most likely to enjoy. By separating the behavior policy from the target policy, Netflix can experiment with different recommendation strategies without compromising the quality of the final recommendations. This approach allows the recommendation system to be both exploratory in gathering new data and exploitative in providing the best possible content to users.
The decoupling of the behavior and target policies allows off-policy methods to reuse experiences more effectively. For instance, experiences collected using a behavior policy that explores the environment broadly can be used to improve the target policy, which aims to maximize rewards. This characteristic makes off-policy methods particularly powerful in dynamic and complex environments where extensive exploration is required [dayan1992q, hessel2018rainbow].
6.3.3 Distinguishing On-Policy vs. Off-Policy Methods
The relationship between the update policy and the behavior policy determines if a method is on-policy or off-policy. Identical policies indicate on-policy, while differing policies indicate off-policy. Implementation details and objectives also influence classification. To better distinguish these methods, we have to first learn what are the different policies.
Behavior Policy is a strategy used by an agent to determine which actions to take at each time step. The behavior policy might, for example, include recommending a variety of movies in order to explore user preferences in the recommendation system example.
Update policy governs how the agent updates its value estimates in response to observed outcomes. Depending on the feedback received from the recommended movies, the update policy of the recommendation system may update the estimated user preferences.A thorough understanding of the interactions between these policies is essential for the implementation of effective learning systems. An agent’s behavior policy determines how it explores an environment, balancing exploration with exploitation to gather useful information. Alternatively, the update policy determines how the agent learns from these experiences in order to improve its estimates of value.
When using on-policy methods, the behavior policy and the update policy are the same, meaning that the actions taken to interact with the environment are also used to update the value estimates. The result is stable learning, but it can be less efficient because the policy may not sufficiently explore the state space [rummery1994line].
There is a difference between the behavior policy and the update policy in off-policy methods. As opposed to the behavior policy, the update policy focuses on optimizing the value estimates by taking the most appropriate action. Despite the fact that this separation can make learning more efficient, it can also introduce instability if the behavior policy diverges too far from the optimal policy [mnih2015human, hessel2018rainbow]. Furthermore, advanced methods, such as Actor-Critic algorithms, separate the behavior policy (actor) and the update policy (critic). Actors make decisions according to current policies, while critics evaluate these decisions and provide feedback to improve policies, thus combining the stability of on-policy methods with the efficiency of off-policy methods [konda1999actor, lillicrap2015continuous].
7 A Timeline of Reinforcement Learning Evolution
In recent decades, reinforcement learning has undergone significant changes. The timeline below highlights key milestones and advancements in the field, beginning with the earliest contributions and moving on to the most recent trends. In addition, the timeline provides readers with an opportunity to gain insight into the developments and potential directions for research.
  • 1950s-1960s: Early Foundations
    • 1954: Marvin Minsky’s SNARC, an early neural network-based machine, was one of the first to use reinforcement principles [minsky1954theory].
    • 1954: Richard Bellman’s work on dynamic programming laid the foundation for modern reinforcement learning [bellman1954theory].
    • 1965: Bellman and Kalaba introduced the Bellman equation, fundamental to many RL algorithms [bellman1965dynamic].
  • 1970s-1980s: Theoretical Development
    • 1972: Klopf’s work on drive-reduction theory in neural networks contributed to the theoretical underpinnings of RL [klopf1972brain].
    • 1984: Barto, Sutton, and Anderson developed the concept of temporal difference (TD) learning, a cornerstone of RL [sutton1984temporal].
    • 1989: Watkins introduced Q-learning, a pivotal off-policy RL algorithm [watkins1989learning].
    • 1989: Sutton and Barto’s book, ”Reinforcement Learning: An Introduction,” became a foundational text in the field [sutton1988learning].
  • 1990s: Early Applications and Further Theory
    • 1992: Watkins and Dayan formalized Q-learning and provided theoretical convergence guarantees [dayan1992q].
    • 1994: Rummery and Niranjan developed SARSA, an on-policy RL algorithm [rummery1994line].
    • 1996: Chris Watkins’ Dyna-Q algorithm combined model-free and model-based approaches [sutton1990integrated].
  • 1999 - 2000s: Advances in Algorithms and Applications
    • 1999: Konda and Tsitsiklis introduced actor-critic methods, blending value and policy-based approaches [konda1999actor].
    • 1999: Ng, Harada, and Russell’s work on apprenticeship learning advanced RL applications in robotics [ng1999policy].
    • 2001: Tsitsiklis and Van Roy developed the theory of least-squares temporal difference (LSTD) learning [tsitsiklis2001regression].
  • 2010s: Deep Reinforcement Learning and Breakthroughs
    • 2013: Mnih et al. introduced Deep Q-Networks (DQN), integrating deep learning with RL, and demonstrated success on Atari games [mnih2013playing].
    • 2016: Silver et al. developed AlphaGo, combining deep RL and Monte Carlo Tree Search, achieving significant success in the game of Go [silver2016mastering].
    • 2017: Schulman et al. proposed Proximal Policy Optimization (PPO), a robust policy gradient method for RL [schulman2017proximal].
    • 2018: Hessel et al. presented Rainbow, combining several DQN improvements into a single algorithm [hessel2018rainbow].
  • Current Trends and Future Directions
    • 2017: Jaderberg et al. introduced ”Population Based Training,” enhancing the efficiency of RL training processes [jaderberg2017population].
    • 2018 Espeholt et al. proposed ”IMPALA” (Importance Weighted Actor-Learner Architectures), addressing scalability in RL [espeholt2018impala].
    • 2019: OpenAI’s Dactyl demonstrated RL’s potential in robotic manipulation by solving a Rubik’s cube with a robotic hand [akkaya2019solving].
    • The recent trends are focusing more on Interpretability under the umbrella of Interpretable and Explainable RL, Hierarchical and Multi-Agent RL, and Integrations with recent advancements such as Transformers, Large Language Models (LLMs), Human-in-the-loop RL, Social Reinforcement Learning, Transfer Learning, Lifelong Learning, and Continual Learning. It can be seen in details in Table 1.
Trends Reference Paper(s)
Interpretability and Explainable RL [madumal2020explainable], [mott2019towards]
Hierarchical RL [bacon2017option], [vezhnevets2017feudal]
Multi-Agent RL [lowe2017multi], [foerster2016learning]
Integrations with Transformers and LLMs [lu2022frozen], [huang2022language]
Human-in-the-loop RL [christiano2017deep], [macglashan2017interactive]
Social Reinforcement Learning [carroll2019utility], [leurent2019social]
Transfer Learning [rusu2016progressive], [wang2016learning]
Lifelong Learning and Continual Learning [khetarpal2022towards], [yoon2017lifelong]
Table 1: Recent trends in reinforcement learning and key reference papers
8 Essential Algorithms: An Overview
The purpose of this section is to provide a brief overview of essential algorithms based on the material discussed so far. In addition, reference papers are provided for each algorithm to facilitate the search for the original source and facilitate the learning process. In addition to providing a brief description of each algorithm, real-world examples are provided as well. In order to obtain detailed information about each algorithm, readers should refer to the reference papers. Table 2 summarizes all the information.
Algorithm Description Type Policy Reference
TD Learning A fundamental method where updates are made based on the difference between predicted and actual rewards. Model-free On-policy [sutton1988learning]
Q-Learning An off-policy algorithm that learns the value of an optimal policy independently of the agent’s actions. Model-free Off-policy [dayan1992q]
SARSA An on-policy algorithm where the policy is updated based on the action taken by the current policy. Model-free On-policy [rummery1994line]
REINFORCE An on-policy algorithm that uses Monte Carlo methods to update the policy. Model-free On-policy [williams1992simple]
Actor-Critic Combines value function (critic) and policy (actor) updates. Model-free On-policy/Off-policy [konda1999actor]
Dyna-Q Combines model-free and model-based methods by integrating planning, acting, and learning. Model-based Off-policy [sutton1990integrated]
DQN Combines Q-learning with deep neural networks to handle high-dimensional state spaces. Model-free Off-policy [mnih2015human]
TRPO Ensures large updates do not destroy the learned policy by enforcing a trust region. Model-free On-policy [schulman2015trust]
PPO Improves on TRPO by simplifying the algorithm while retaining its performance. Model-free On-policy [schulman2017proximal]
SAC An off-policy actor-critic algorithm that maximizes both expected reward and entropy. Model-free Off-policy [haarnoja2018soft]
Table 2: Essential Reinforcement Learning Algorithms
8.1 TD Learning
Temporal Difference (TD) Learning is a fundamental method in reinforcement learning where updates are made based on the difference between predicted and actual rewards. This approach combines ideas from Monte Carlo methods and dynamic programming [sutton1988learning].
The update rule for TD Learning is:
$V(s)\leftarrow V(s)+\alpha\left[R_{t+1}+\gamma V(S_{t+1})-V(s)\right]$ (37)
where $V(s)$ is the value of state $s$, $\alpha$ is the learning rate, $R_{t+1}$ is the reward received after taking an action, $\gamma$ is the discount factor, and $S_{t+1}$ is the next state.
In stock trading, TD Learning can be used to predict future stock prices based on past price movements. By continuously updating the predicted value of each state (price), traders can make more informed decisions on buying or selling stocks.
8.2 Q-Learning
Q-Learning is an off-policy algorithm that learns the value of an optimal policy independently of the agent’s actions. It aims to find the best action to take in each state to maximize the total reward [dayan1992q].
The update rule for Q-Learning is:
$Q(s,a)\leftarrow Q(s,a)+\alpha\left[r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})-Q(s,a)\right]$ (38)
where $Q(s,a)$ is the value of taking action $a$ in state $s$, $\alpha$ is the learning rate, $r$ is the reward received, $\gamma$ is the discount factor, $s^{\prime}$ is the next state, and $a^{\prime}$ is the next action.
In robotics, Q-Learning can be used to train a robot to navigate through a maze. The robot learns the optimal path to the exit by updating the values of actions taken in each state, even when it explores suboptimal paths during training.
8.3 State-Action-Reward-State-Action (SARSA)
SARSA is an on-policy algorithm where the policy is updated based on the action taken by the current policy. It stands for the sequence of states and actions used to update the Q-values [rummery1994line].
The update rule for SARSA is:
$Q(s,a)\leftarrow Q(s,a)+\alpha\left[r+\gamma Q(s^{\prime},a^{\prime})-Q(s,a)\right]$ (39)
where $Q(s,a)$ is the value of taking action $a$ in state $s$, $\alpha$ is the learning rate, $r$ is the reward received, $\gamma$ is the discount factor, $s^{\prime}$ is the next state, and $a^{\prime}$ is the next action taken according to the policy.
In video games, SARSA can be used to train game agents to make decisions based on the current game state and the actions actually taken, ensuring that the learning process aligns with the agent’s behavior, leading to stable and consistent performance.
8.4 REINFORCE
REINFORCE is an on-policy algorithm that uses Monte Carlo methods to update the policy. It directly optimizes the policy by following the gradient of expected rewards [williams1992simple].
The update rule for REINFORCE is:
$\theta\leftarrow\theta+\alpha\nabla_{\theta}\log\pi_{\theta}(a|s)G_{t}$ (40)
where $\theta$ are the policy parameters, $\alpha$ is the learning rate, $\pi_{\theta}(a|s)$ is the policy probability of taking action $a$ in state $s$, and $G_{t}$ is the return (total accumulated reward).
In personalized marketing, REINFORCE can be used to adjust marketing strategies based on user interactions and feedback. By optimizing the policy to maximize customer engagement, businesses can improve their marketing effectiveness.
8.5 Actor-Critic
Actor-Critic methods combine value function (critic) and policy (actor) updates. The actor updates the policy parameters in the direction suggested by the critic, which evaluates the action-value function [konda1999actor].
The update rules for Actor-Critic are:
$\delta_{t}=R_{t+1}+\gamma V(S_{t+1})-V(S_{t})$ (41)
$\theta\leftarrow\theta+\alpha\delta_{t}\nabla_{\theta}\log\pi_{\theta}(A_{t}|S_{t})$ (42)
where $\delta_{t}$ is the temporal difference error, $\theta$ are the policy parameters, $\alpha$ is the learning rate, $\pi_{\theta}(A_{t}|S_{t})$ is the policy probability, and $V(S_{t})$ is the value function.
In autonomous driving, Actor-Critic methods can be used to improve the driving policy by having the actor suggest actions and the critic evaluate the consequences. This leads to safer and more efficient driving behaviors.
8.6 Soft Actor-Critic
Soft Actor-Critic (SAC) is an off-policy actor-critic algorithm that maximizes both expected reward and entropy. It aims to encourage exploration by adding an entropy term to the objective [haarnoja2018soft].
The update rules for SAC are:
$\mathcal{L}(\theta)=\mathbb{E}_{(s,a)\sim D}\left[-Q(s,a)+\alpha\log\pi_{\theta}(a|s)\right]$ (43)
where $\alpha$ is the temperature parameter controlling the trade-off between exploration and exploitation.
In energy management systems, SAC can be used to optimize energy consumption by balancing the trade-off between energy savings and operational efficiency, ensuring robust and adaptable policies.
8.7 Dyna-Q
Dyna-Q combines model-free and model-based methods by integrating planning, acting, and learning. It uses a learned model of the environment to simulate experiences and update the policy [sutton1990integrated].
The update rule for Dyna-Q is:
$Q(s,a)\leftarrow Q(s,a)+\alpha\left[r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})-Q(s,a)\right]$ (44)
and additionally, simulated experiences are used to update the Q-values.
In logistics, Dyna-Q can be used to optimize delivery routes by simulating different scenarios and learning from both real and simulated experiences. This improves route planning and efficiency.
8.8 Deep Q-Network
Deep Q-Network (DQN) combines Q-learning with deep neural networks to handle high-dimensional state spaces. It uses a neural network to approximate the Q-values [mnih2015human].
The update rule for DQN is similar to Q-learning, but the Q-values are approximated using a neural network:
$Q(s,a;\theta)\leftarrow Q(s,a;\theta)+\alpha\left[r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime};\theta^{\prime})-Q(s,a;\theta)\right]$ (45)
where $\theta$ are the neural network parameters.
In playing complex video games like Atari, DQN can be used to train agents to play the games at superhuman levels by learning directly from pixel inputs and using neural networks to approximate the optimal policies.
8.9 Trust Region Policy Optimization
Trust Region Policy Optimization (TRPO) ensures large updates do not destroy the learned policy by enforcing a trust region. It uses a constraint to limit the step size of policy updates [schulman2015trust].
The update rule for TRPO is:
$\max_{\theta}\mathbb{E}_{s\sim\rho^{\pi},a\sim\pi_{\theta}}\left[\frac{\pi_{\theta}(a|s)}{\pi_{\theta_{\text{old}}}(a|s)}A^{\pi_{\theta_{\text{old}}}}(s,a)\right]$ (46)
subject to:
$\mathbb{E}_{s\sim\rho^{\pi}}\left[D_{KL}\left(\pi_{\theta_{\text{old}}}(\cdot|s)||\pi_{\theta}(\cdot|s)\right)\right]\leq\delta$ (47)
where $D_{KL}$ is the Kullback-Leibler divergence and $\delta$ is a small positive number.
In robotic control, TRPO can be used to optimize the movements of robotic arms, ensuring stable and efficient learning of complex tasks without drastic policy changes.
8.10 Proximal Policy Optimization
Proximal Policy Optimization (PPO) improves on TRPO by simplifying the algorithm while retaining its performance. It uses a clipped objective to limit policy updates [schulman2017proximal].
The update rule for PPO is:
$\mathcal{L}^{\text{CLIP}}(\theta)=\mathbb{E}_{t}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\text{clip}\left(r_{t}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{t}\right)\right]$ (48)
where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$, $\hat{A}_{t}$ is the advantage estimate, and $\epsilon$ is a small positive number.
In virtual personal assistants, PPO can be used to train the assistant to handle various tasks efficiently while ensuring stable improvements in performance.
9 Resources to Learn Reinforcement Learning
It is a complex set of problems and applications within RL that require a greater level of research resources. By making these resources available, it is not only possible to gain a deeper understanding of theoretical concepts and practical tools, but it will also contribute to the advancement of state-of-the-art research through the support of the community. The importance of comprehensive resources for successful real-life research is highlighted in this chapter, and we provide a list of some of the most helpful books, courses, videos, and online communities to assist readers in getting started with their real-life research without being overwhelmed.
9.1 Books
We provide a list of highly recommended books which provide foundational knowledge and practical insight into reinforcement learning. A comprehensive overview of basic concepts, advanced algorithms, and practical examples is provided in these books, which will provide you with an understanding of the field on a deeper level.
  1. 1.
    ”Reinforcement Learning: An Introduction” by Richard S. Sutton and Andrew G. Barto
    • The foundational book on reinforcement learning, covering basic concepts, algorithms, and theoretical foundations.
  2. 2.
    ”Deep Reinforcement Learning Hands-On” by Maxim Lapan
    • Practical guide to deep reinforcement learning using Python and PyTorch, with hands-on examples.
  3. 3.
    ”Grokking Deep Reinforcement Learning” by Miguel Morales
    • A beginner-friendly introduction to deep reinforcement learning with intuitive explanations and hands-on exercises.
  4. 4.
    ”Algorithms for Reinforcement Learning” by Csaba Szepesvári
    • A concise and clear introduction to reinforcement learning algorithms.
9.2 Online Courses
The following is a list of online courses that offer comprehensive reinforcement learning experiences. There is a range of courses available, from introductory to advanced levels, covering fundamental concepts and practical implementations along with real-life applications.
  1. 1.
    Coursera: Reinforcement Learning Specialization by University of Alberta
    • A comprehensive series of courses covering fundamental concepts, algorithms, and applications of reinforcement learning.
  2. 2.
    Udacity: Deep Reinforcement Learning Nanodegree
    • Focuses on deep reinforcement learning using PyTorch, including projects and practical implementations.
  3. 3.
    edX: Fundamentals of Reinforcement Learning by University of Alberta
    • An introduction to the fundamentals of reinforcement learning, including key algorithms and techniques.
  4. 4.
    Reinforcement Learning Winter 2019
    • Lecture videos and materials from Stanford’s advanced reinforcement learning course.
9.3 Video Lectures
The following are links to video lecture series produced by leading researchers and institutions. A variety of topics in RL are covered in these lectures, providing valuable insights and knowledge from experts in the field.
  1. 1.
    DeepMind x UCL — Reinforcement Learning Lectures Series
    • A series of lectures by leading researchers in reinforcement learning, covering fundamental and advanced topics.
  2. 2.
    David Silver’s Reinforcement Learning Course
    • Lectures by David Silver, a principal researcher at DeepMind, providing an in-depth introduction to reinforcement learning.
  3. 3.
    Pascal Poupart’s Reinforcement Learning Course - CS885
    • Lectures by Pascal Poupart, an esteemed professor at the University of Waterloo, providing an in-depth introduction to reinforcement learning.
  4. 4.
    Sarath Chandar’s Reinforcement Learning Course
    • Lectures by Sarath Chandar, principal investigator of the Chandar Research Lab (CRL) in the Department of Computer and Software Engineering at Polytechnique Montreal.
9.4 Tutorials and Articles
These resources include educational materials, tutorials, and articles that provide practical guidance on implementing RL algorithms. Code examples are provided as well as theoretical background and step-by-step instructions for various RL techniques in these resources.
  1. 1.
    OpenAI Spinning Up in Deep RL
    • An educational resource from OpenAI that provides an introduction to deep reinforcement learning, including code examples and theoretical background.
    • Link: Spinning Up
  2. 2.
    Deep Reinforcement Learning Course by PyTorch
    • A free online course provided by PyTorch, focusing on practical implementations of deep reinforcement learning.
  3. 3.
    RL Adventure by Denny Britz
[…]

### v2

@@ 3.2 Off-Policy and On-Policy Methods
Off-policy methods, on the other hand, involve learning the value of the optimal policy independently of the agent’s actions. In these methods, we distinguish between two types of policies: the behavior policy ($b$) and the target policy ($\pi$). The behavior policy explores the environment, while the target policy aims to improve performance based on the gathered experience. This allows for a more exploratory behavior policy while learning an optimal target policy. A significant advantage of off-policy methods is that they can learn from data generated by any policy, not just the one currently being followed, making them highly flexible and sample-efficient. The decoupling of the behavior and target policies allows off-policy methods to reuse experiences more effectively. For instance, experiences collected using a behavior policy that explores the environment broadly can be used to improve the target policy, which aims to maximize rewards. This characteristic makes off-policy methods particularly powerful in dynamic and complex environments where extensive exploration is required [24, 25].
The relationship between the target policy and the behavior policy determines if a method is on-policy or off-policy. Identical policies indicate on-policy, while differing policies indicate off-policy. Implementation details and objectives also influence classification. To better distinguish these methods, we have to first learn what are the different policies. Behavior Policy $b$ is a strategy used by an agent to determine which actions to take at each time step. The behavior policy might, for example, include recommending a variety of movies in order to explore user preferences in the recommendation system example. Target policy $\pi$ governs how the agent updates its value estimates in response to observed outcomes. Depending on the feedback received from the recommended movies, the target policy of the recommendation system may update the estimated user preferences.A thorough understanding of the interactions between these policies is essential for the implementation of effective learning systems. An agent’s behavior policy determines how it explores an environment, balancing exploration with exploitation to gather useful information. Alternatively, the target policy determines how the agent learns from these experiences in order to improve its estimates of value. When using on-policy methods, the behavior policy and the target policy are the same, meaning that the actions taken to interact with the environment are also used to update the value estimates. The result is stable learning, but it can be less efficient because the policy may not sufficiently explore the state space [26]. There is a difference between the behavior policy and the target policy in off-policy methods. As opposed to the behavior policy, the target policy focuses on optimizing the value estimates by taking the most appropriate action. Despite the fact that this separation can make learning more efficient, it can also introduce instability if the behavior policy diverges too far from the optimal policy [15, 25]. Furthermore, advanced methods, such as Actor-Critic algorithms, separate the behavior policy (actor) and the target policy (critic). Actors make decisions according to current policies, while critics evaluate these decisions and provide feedback to improve policies, thus combining the stability of on-policy methods with the efficiency of off-policy methods [27, 28].
Understanding the core methodologies of RL, such as model-free and model-based approaches, as well as the distinction between off-policy and on-policy methods, provides a foundational framework for exploring the diverse landscape of RL algorithms. These methodologies not only shape how agents learn from their environments but also influence their adaptability and efficiency in complex, real-world scenarios. Building on this understanding, the next section delves deeper into the specific essential algorithms underpinning these methods, focusing on the policy-based, value-based, and hybrid approaches, to provide a clearer picture of their mechanisms and applications in RL.
4 Essential Algorithms
This section aims to present a concise overview of key algorithms discussed thus far, accompanied by references to the original research papers for further exploration. Each algorithm is briefly described, and real-world examples are included to enhance understanding. Readers seeking detailed information are encouraged to consult the cited references, which serve as a gateway to the primary sources and support a deeper learning experience. We categorize algorithms into three types: Value-based, Policy-based, and Hybrid Algorithms. For each type, we analyze one to two widely-used algorithms, acknowledging that there are more algorithms to discover.
4.1 Value-based
We introduced value-based methods, and clearly analyzed how they work. Here, we introduce three algorithms, that are in the tabular settings. Later, we dive deeper into the topics, and discuss value-based methods that use Deep Learning, such as Deep Q-Networks.
A significant breakthrough was made by [29] with the introduction of Q-learning, a Model-free algorithm considered as off-policy Temporal Difference (TD) control. TD learning is undoubtedly the most fundamental and innovative concept. A combination of Monte Carlo (MC) methods and Dynamic Programming (DP) is used in this method. On one hand, similar to MC approaches, TD learning can be used to acquire knowledge from unprocessed experience without the need for a model that describes the dynamics of the environment. On the other hand, TD algorithms are also similar to DP in that they refine predictions using previously learned estimates instead of requiring a definitive outcome in order to proceed (known as bootstrapping). Q-learning enables an agent to learn the value of an action in a particular state through experience, without requiring a model of the environment. It operates on the principle of learning an action-value function that gives the expected utility of taking a given action in each state and following a fixed policy thereafter. The core of the Q-learning algorithm involves updating the Q-values (action-value pairs), where the learned action-value function, denoted as $\mathcal{Q}$, approximates $q_{*}$, the optimal action-value function, regardless of the policy being followed. This significantly simplifies the algorithm’s analysis and has facilitated early proofs of convergence. However, the policy still influences the process by determining which state-action pairs are visited and subsequently updated.
$\mathcal{Q}(\mathcal{S}_{t},\mathcal{A}_{t})\leftarrow\mathcal{Q}(\mathcal{S}_{t},\mathcal{A}_{t})+\alpha\big{[}\mathcal{R}_{t+1}+\gamma\max_{a}\mathcal{Q}(\mathcal{S}_{t+1},a)-\mathcal{Q}(\mathcal{S}_{t},\mathcal{A}_{t})\big{]}$ (34)
There are other types of Q-learning introduced in the literature, with slight changes and improvements, such as: Double Q-learning [30], that addresses the overestimation bias in Q-learning, Distributional Q-learning [31], which models the distribution of returns instead of estimating the mean Q-value, providing richer information for decision-making, and many more [32, 33, 34].
Another widely used value-based algorithm is Deep Q-Networks (DQN), which merges Q-learning with Neural Networks to learn control policies directly from raw pixel inputs. It uses Convolutional Neural Networks (CNN) to process these inputs and an experience replay mechanism to stabilize learning by breaking correlations between consecutive experiences. The target network, updated less frequently, aids in stabilizing training. DQN achieved state-of-the-art performance on various Atari 2600 games, surpassing previous methods and, in some cases, human experts, using a consistent network architecture and hyperparameters across different games [17]. DQN combines the introduced Bellman Equation with DL approaches like Loss Function and Gradient Descent to find the optimal policy as below:
$\mathcal{L}_{i}(\theta_{i})=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\left[\left(y_{i}-\mathcal{Q}(s,a;\theta_{i})\right)^{2}\right]$ (35)
where
$y_{i}=r+\gamma\max_{a^{\prime}}\mathcal{Q}(s^{\prime},a^{\prime};\theta^{-})$ (36)
$\nabla_{\theta_{i}}\mathcal{L}_{i}(\theta_{i})=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\Big{[}$ $\left(r+\gamma\max_{a^{\prime}}\mathcal{Q}(s^{\prime},a^{\prime};\theta^{-})\right.\left.-\mathcal{Q}(s,a;\theta_{i})\right)\nabla_{\theta_{i}}\mathcal{Q}(s,a;\theta_{i})\Big{]}$ (37)
Similar to Q-learning, there have been updates made to DQN. Some of the variations are [35, 36, 37].
Table 1: Essential RL Algorithms
Algorithms
1 Q-Learning [29] - Model-free, Off-policy, Value-based
2 SARSA (State-Action-Reward-State-Action) [26] - Model-free, On-policy, Value-based
3 Expected SARSA [38] - Model-free, On-policy, Value-based
4 REINFORCE [39] - Model-free, On-policy, Policy-based
5 Dyna-Q [40] - Model-based, Off-policy, Hybrid
6 DQN [17] - Model-free, Off-policy, Value-based
7 TRPO [41] - Model-free, On-policy, Policy-based
8 PPO [18] - Model-free, On-policy, Policy-based
9 SAC (Soft Actor-Critic) [33] - Model-free, Off-policy, Hybrid
10 A3C [42] - Model-free, On-policy, Hybrid
11 A2C [28] - Model-free, On-policy, Hybrid
12 DDPG (Deep Deterministic Policy Gradient) [28] - Model-free, Off-policy, Policy-based
13 TD3 (Twin Delayed Deep Deterministic Policy Gradient) [43] - Model-free, Off-policy, Policy-based
4.2 Policy-based
Moving on from value-based methods, we analyze some policy-based algorithms in this section. Policy-based methods are another fundamental RL method that more strongly emphasizes direct policy optimization in the process of choosing actions for an agent. In contrast to Value-based methods, which search for the value function implicit in the task, and then derive an optimal policy, Policy-based methods directly parameterize and optimize the policy. This approach offers several advantages, particularly better dealing with very challenging environments that have high-dimensional action spaces or where policies are inherently stochastic. Perhaps at the core, Policy-based methods conduct their operation based on the parameterization of policies, usually denoted as $\pi(a|s;\theta)$. Here, $\theta$ is used to denote the parameters of the policy, while $s$ denotes the state and $a$ denotes the action. In other words, it finds the optimal parameters $\theta^{*}$ that maximize the expected cumulative reward. Needless to say, this is generally done by gradient ascent techniques and more specifically by Policy Gradient methods that explicitly compute the gradient of expected reward with respect to the policy parameters, modifying parameters in the direction of reward increase [1, 10, 44].
REINFORCE is one of the widely-used policy-based algorithms. The REINFORCE algorithm is a seminal contribution to RL, particularly within the context of policy gradient methods. The algorithm is designed to optimize the expected cumulative reward by adjusting the policy parameters in the direction of the gradient of the expected reward. It is rooted in the stochastic policy framework, where the policy, parameterized by $\theta$, defines a probability distribution over actions given the current state. The key insight of the REINFORCE algorithm is to use the log-likelihood gradient estimator to update the policy parameters [39]. The gradient of the expected reward with respect to the policy parameters $\theta$ is given by:
$\nabla_{\theta}\mathcal{J}(\theta)=\mathbb{E}_{\pi}\left[\nabla_{\theta}\log\pi_{\theta}(a|s)\mathcal{G}_{t}\right],$ (38)
where $\pi_{\theta}(a|s)$ is the probability of taking action $a$ in the state $s$ under policy $\pi$ parameterized by $\theta$, and $\mathcal{G}_{t}$ is the return (cumulative future reward) following time step $t$. This gradient estimation forms the basis for the parameter update rule.
$\theta\leftarrow\theta+\alpha\nabla_{\theta}\log\pi_{\theta}(a|s)\mathcal{G}_{t},$ (39)
where $\alpha$ is the learning rate.
Another algorithm, which has been used in variety of applications [45, 46, 47, 48, 49, 50, 51, 52], is Proximal Policy Optimization (PPO). PPO, proposed by [18], represents a significant advancement within policy gradient methods. PPO aims to achieve reliable performance and sample efficiency, addressing the limitations of previous policy optimization algorithms such as Vanilla Policy Gradient (VPG) [53] methods and Trust Region Policy Optimization (TRPO) [41]. Using policy gradient methods, the policy parameters are optimized through stochastic gradient ascent by estimating the gradient of the policy. One of the most commonly used policy gradient estimators is:
$\hat{g}=\hat{\mathbb{E}}_{t}\left[\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})\hat{A}_{t}\right],$ (40)
where $\pi_{\theta}$ represents the policy parameterized by $\theta$, and $\hat{A}_{t}$ is an estimator of the advantage function at time step $t$. This estimator helps construct an objective function whose gradient corresponds to the policy gradient estimator:
$\mathcal{L}_{\mathcal{PG}}(\theta)=\hat{\mathbb{E}}_{t}\left[\log\pi_{\theta}(a_{t}|s_{t})\hat{A}_{t}\right].$ (41)
PPO simplifies TRPO by using a surrogate objective with a clipped probability ratio, allowing for multiple epochs of mini-batch updates. In order to preserve learning, large policy updates should be avoided.
4.3 Hybrid (Actor-Critic) methods
For the last group of algorithm, hybrid methods, we introduce Asynchronous Advantage Actor-Critic (A3C) and Advantage Actor-Critic (A2C). Actor-critic methods combine Value-based and Policy-based approaches. Essentially, these methods consist of two components: the Actor, who selects actions based on a policy, and the Critic, who evaluates the actions based on their value function. By providing feedback on the quality of the actions taken, the critic guides the actor in updating the policy directly. As a result of this synergy, learning can be more stable and efficient, addressing some limitations of pure policy or Value-based approaches [16, 54]. The A2C algorithm is a synchronous variant of the A3C algorithm, which was introduced by [42]. A2C maintains the key principles of A3C but simplifies the training process by synchronizing the updates of multiple agents, thereby leveraging the strengths of both Actor-Critic methods and advantage estimation. The Actor-Critic architecture combines two primary components, in both algorithms: the actor, which is responsible for selecting actions, and the critic, which evaluates the actions by estimating the value function. The actor updates the policy parameters in a direction that is expected to increase the expected reward, while the critic provides feedback by computing the TD error. This integration allows for more stable and efficient learning compared to using Actor-only or critic-only methods [27]. Advantage estimation is a technique used to reduce the variance of the policy gradient updates. The advantage function $\mathcal{A}(s,a)$ represents the difference between the action-value function $\mathcal{Q}(s,a)$ and the value function $\mathcal{V}(s)$.
$\mathcal{A}(s,a)=\mathcal{Q}(s,a)-\mathcal{V}(s).$ (42)
By using the advantage function, A2C focuses on actions that yield higher returns than the average, which helps in making more informed updates to the policy [1]. Unlike A3C, where multiple agents update the global model asynchronously, A2C synchronizes these updates. Multiple agents run in parallel environments, collecting experiences and calculating gradients, which are then aggregated and used to update the global model synchronously. This synchronization reduces the complexity of implementation and avoids issues related to asynchronous updates, such as non-deterministic behavior and potential overwriting of gradients.
Table 1 categorizes the examined algorithms, and some other algorithms that we did not discuss to give a comprehensive overview in terms of
Table 2: Reinforcement Learning Resources
Books
1 ”Reinforcement Learning: An Introduction” by Richard S. Sutton and Andrew G. Barto
2 ”Deep Reinforcement Learning Hands-On” by Maxim Lapan
3 ”Grokking Deep Reinforcement Learning” by Miguel Morales
4 ”Algorithms for Reinforcement Learning” by Csaba Szepesvári
Online Courses
5 Coursera: Reinforcement Learning Specialization by University of Alberta
6 Udacity: Deep Reinforcement Learning Nanodegree
7 edX: Fundamentals of Reinforcement Learning by University of Alberta
8 Reinforcement Learning Winter 2019 (Stanford)
Video Lectures
9 DeepMind x UCL — Reinforcement Learning Lecture Series
10 David Silver’s Reinforcement Learning Course
11 Pascal Poupart’s Reinforcement Learning Course - CS885
12 Sarath Chandar’s Reinforcement Learning Course
Tutorials and Articles
13 OpenAI Spinning Up in Deep RL
14 Deep Reinforcement Learning Course by PyTorch
15 RL Adventure by Denny Britz
Online Communities and Forums
16 Reddit: r/reinforcementlearning
17 Stack Overflow
18 AI Alignment Forum
[…]